OpenStudy (anonymous):

(4+5i)(5+i) multiply

OpenStudy (abb0t):

Are the i's imaginary numbers or just regular i's?

OpenStudy (anonymous):

Use the FOIL method- multiply the first 2 terms, then the outer two, then inner two, and then the last ones. Remember that if you have an i^2, it's a -1. Ex: (a+b)(c+d) ac+ad+cb+bd

OpenStudy (anonymous):

NOTE: i^2=-1 (4+5i)(5+i) (4*5i)+(4*5)+(5i*5)+(5i*i) 20i+20+25i+5i^2 20+45i+5i^2 and i^2=-1 so 20+45i+5*(-1) 45i+20-5 45i+15 should be your answer